Aircraft arrive at the Mckinnon airport with a poisson distribution at a rate of one aircraft arrival every 35 minutes. What is the chance that at least 2 aircraft arrive in a 20-minute window of time?
Answer #1

arrival rate of aircraft = 60/35 = 12/7 per hour

or 12/7 x1/3 per 20 min = 4/7 per 20 min.

Probability that at least 2 aircraft will arrive

1- P(r<2) = 1- sum ( r =0 to1) exp (-4/7) (4/7)r / factorial r

= 1- exp(-4/7) [ 1+4/7]

= 1- 0.5647 x 1.5714

= 1-0.8873 =0.1126